OpenStudy (anonymous):

write the equation passing through (5,7) and parallel to 7x-8y=10? please..

OpenStudy (anonymous):

In slope intercept form, or point slope form?>

OpenStudy (anonymous):

slope intercept form.

OpenStudy (anonymous):

heres my answer key.... y=-7/8x+21/8... y=7/8x+5...y=7/8x-21/8....y=7/8x+21/8

OpenStudy (anonymous):

i got y=7/8x and then im confused with the answer choices, because i didnt get any of those

OpenStudy (anonymous):

  None of the choices match though.. o.o